# Primordial black hole formation in $k$-inflation models

### v3

@@ 1 Introduction
There are various scenarios for primordial black hole production. We aim to investigate PBH production in the early Universe in two inflation models with noncanonical kinetic terms. The models in which the kinetic term has a noncanonical form are usually referred to as $k$-essence [18] or, in the context of inflation, $k$-inflation models [19]. In this work, we study two $k$-essence models using the Hamiltonian formalism for the background field equations.
First, we investigate a $k$-essence model proposed by Papanikolaou, Lymperis, Lola, and Saridakis in Ref. [20]. We will refer to their model as the PLLS model. The PLLS Lagrangian is of the form
@@
where $\varphi$ is a scalar field and $\alpha\geq 0$ is a constant. The same form of $k$-essence has recently been studied in the context of the black-bounce solution [21]. The potential $U$ in Ref. [20] is assumed to be a positive function of $\varphi$ with a specific inflection point. An inflationary potential with an inflection or near inflection point leads to an ultra slow-roll regime, which in turn leads to a peak in the power spectrum of curvature perturbations on comoving wavenumber scales above the CMB pivot scale $q_{\rm CMB}$. As a result, the enhanced cosmological perturbations could collapse into PBHs [9, 22, 23, 24, 25, 26].
@@
The Tachyon model is motivated by low-energy string theory. The existence of tachyons in the perturbative spectrum of string theory shows that the perturbative vacuum is unstable. This instability implies that there exists a true vacuum to which a tachyon field $\varphi$ tends [45]. The foundations of this process are given by a model of effective field theory [46] with a Lagrangian of the Dirac-Born-Infeld (DBI) form
@@
A similar Lagrangian appears in the so-called brane inflation models [47]. In these models, the motion of a D3-brane in a warped throat region of a compact space drives inflation, and the DBI field corresponds to the position of the D3-brane. Here, we assume that the tachyon potential has an inflection point, as in the PLLS case.
@@
We use the metric signature (+,,,)(+,-,-,-). Unless specified otherwise, we use the system of units in which $c=\hbar=1$. For convenience, we introduce a length scale $\ell$ so that $\ell\gg 1/M_{\rm Pl}$, where $M_{\rm Pl}=\sqrt{1/(8\pi G)}$ is the reduced Planck mass. By $\mathcal{L}$ and $\mathcal{H}$, we denote the Lagrangian and Hamiltonian multiplied by $\ell^{4}$ so that our $\mathcal{L}$ and $\mathcal{H}$ are dimensionless. By $\varphi$ and $\phi$, we denote the scalar fields connected by $\varphi=\ell^{2}\phi$ and have the dimension of length and mass, respectively.
@@ 2 The PLLS Model
where we assume that $X\equiv g^{\mu\nu}\varphi_{,\mu}\varphi_{,\nu}$ is postive. Here, $\varphi$ is a scalar field of dimension of length, and $\alpha\geq 1$ is a constant. The physical field $\phi$, related to $\varphi$ via
@@
The background evolution will be studied using the covariant Hamilton formalism. (for details see Ref. [48]). The Hamiltonian corresponding to a $k$-essence Lagrangian is obtained by the Legendre transformation
${\cal H}=2X{\cal L}_{X}-{\cal L},$ (6)
where the subscript $X$ denotes a derivative with respect to $X$. We find
@@
Now we introduce the variable $\eta$ conjugate to $\sqrt{X}$ as described in [48]
$\eta=\epsilon\frac{\partial{\cal L}}{\partial\sqrt{X}},$ (10)
where $\epsilon$ is +1+1 or 1-1 according to whether $\varphi_{,0}$ is positive or negative. In this model, the variable $X$ can be expressed as an explicit function of $\eta$. By making use of (3) and (10) we find
$\frac{X}{2}=\left(\frac{\eta^{2}}{2\alpha^{2}}\right)^{1/(2\alpha-1)}.$ (11)
@@
The Hamilton equations
$\dot{\varphi}=\frac{\partial{\cal H}}{\partial\eta},$ (15)
$\dot{\eta}+3H\eta=-\frac{\partial{\cal H}}{\partial\varphi}.$ (16)
govern the background dynamics. Introducing the e-fold number
@@
Instead of the two first-order Hamilton equations, one can use the second-order field equation of motion. This equation can be obtained directly from the Lagrangian or combining Eqs. (18) and (19). Either way one finds111 A comparison with Ref. [20] reveals that the factor $\alpha(2\alpha-1)$ in the denominator of the last term on the left-hand side of Eq. (21) is missing in their equation (2.9).
@@ 3 The Tachyon model
As in the PLLS model, we choose the potential $U(\varphi)$ defined by (5). The parameters of the potential are $\ell=10^{6}M_{\rm Pl}^{-1}$, $n=3$, and $V_{0}=10^{-16}M_{\rm Pl}^{4}$ as before. For this model we choose a slightly different $\lambda=7.502\cdot 10^{-6}$. Regarding the initial values for $\varphi$ and $\eta$, we first choose $\phi_{\rm in}$ in units of $M_{\rm Pl}$ of the order of those in Ref. [20] with the corresponding $\varphi_{\rm in}=\phi_{\rm in}/\ell^{2}$. Then, we tune $\varphi_{\rm in}$ and $\eta_{\rm in}$ to obtain plateau-like solutions as shown in Fig. 2.
@@ 4 Curvature perturbations
The power spectrum of the curvature perturbations is given by
@@
where $\zeta_{q}(N_{q})$ is the solution to Eq. (37) taken at the point $N_{q}$ which for given $q$ satisfy (38). The curvature perturbation spectrum needs to satisfy the Harrison-Zeldovich spectrum near $q=q_{\rm CMB}=0.05$ Mpc-1
@@
where the conformal time $\tau=\int dt/a$ in the slow-roll regime satisfies
$\tau=-\frac{1+\varepsilon_{1}}{aH}+\mathcal{O}(\varepsilon_{1}^{2}),$ (50)
with $qc_{\rm s}\tau\gg 1$ in the deep subhorizon region. Hence, the initial values of $v_{q}$ and $v_{q}^{\prime}$ at $N_{{\rm in},q}$ are determined by (49) up to an arbitrary phase. The simplest choice is
@@ 4.2 Approximate spectrum
where for each $q$, the quantities $H$, $c_{\rm s}$, and $\varepsilon_{1}$ take on their horizon crossing values at the corresponding $N_{q}$. The approximate spectrum obtained in this way will not in general have the correct observed value at the pivot scale $q_{\rm CMB}$. To satisfy the proper normalization of $\mathcal{P}_{\rm S}(q)$ at $q=q_{\rm CMB}$, we can introduce a constant factor $\bar{c}_{0}$. To wit, we define
@@ 5 PBH formation
The quantity $\zeta(r)$ is spherically symmetric comoving curvature perturbation conserved on superhorizon scales [53]. The corresponding density contrast is given by [9]
$\delta_{r}=-\frac{8}{9}\frac{1}{a^{2}H^{2}}e^{-5\zeta(r)/2}\nabla^{2}e^{\zeta(r)/2}\simeq-\frac{4}{9}\frac{1}{a^{2}H^{2}}\nabla^{2}\zeta(r),$ (59)
or in the momentum space
@@
The power spectrum of the density contrast is then
@@
We also define the smoothed variance of the density contrast over the smoothing scale $R=1/(aH)$
@@
where $W$ is a smoothing window function
$W(q,R)=e^{-q^{2}R^{2}/2}.$ (64)
@@ 5.1 PBH mass fraction
Press-Schechter formalism
In the Press-Schechter formalism of gravitational collapse, the mass fraction of PBHs of mass $M$ is given by the probability that the overdensity $\delta$ is above a certain threshold value $\delta_{c}$ for collapse. Assuming $\delta$ is a Gaussian random variable with mass (or scale) dependent variance, the mass fraction $\beta(M)$ at the time of formation is then given by [56, 57, 58]
@@
Critical collapse and peaks (CCP) formalism
@@
The value of the critical collapse threshold $\delta_{c}$ depends on the shape of the collapsing curvature power spectrum [63] (see also [57, 64]). In appendix A, we outline the calculation procedure. For a broad power spectrum that could be approximated by (65), one finds $\delta_{c}\sim 0.56$ [63].
Comparison
To compare these two methods, we can use the toy-model spectrum (65). In this model, the function $\beta(M)$ may be approximated by a similar top-hat profile

### v4

@@ 1 Introduction
There are various scenarios for primordial black hole production, such as: formation of PBHs during the reheating phase at the end of inflation [18], inflaton potential with inflection [9, 19, 20, 21, 22, 23, 24, 25], sound speed resonace during inflation [26, 27], null-energy condition violation during inflation [28], first order phase transitions [29], and inflation models with non-canonical kinetic terms [30, 31].
We aim to investigate PBH production in the early Universe in two inflation models both with a potential with inflection and noncanonical kinetic terms, in particular the Dirac-Born-Infeld (DBI) type of Lagrangian. The models in which the kinetic term has a noncanonical form are usually referred to as $k$-essence [32] or, in the context of inflation, $k$-inflation models [33]. In this work, we study two $k$-essence models using the Hamiltonian formalism for the background field equations.
First, we investigate a $k$-essence model proposed by Papanikolaou, Lymperis, Lola, and Saridakis [31]. We will refer to their model as the PLLS model. The reason why we study this model is twofold: a) the model serves as a good example worth comparison with another model we study in this paper, and b) the results obtained here are modified due to correcting a numerical mistake of Ref. [31](see footnote 1 on page 6). The PLLS Lagrangian is of the form
@@
where $\varphi$ is a scalar field and $\alpha\geq 0$ is a constant. The same form of $k$-essence has recently been studied in the context of the black-bounce solution [34]. The potential $U$ in Ref. [31] is assumed to be a positive function of $\varphi$ with a specific inflection point. An inflationary potential with an inflection or near inflection point leads to an ultra slow-roll regime, which in turn leads to a peak in the power spectrum of curvature perturbations on comoving wavenumber scales above the CMB pivot scale $q_{\rm CMB}$. As a result, the enhanced cosmological perturbations could collapse into PBHs.
@@
The Tachyon model is motivated by low-energy string theory. The existence of tachyons in the perturbative spectrum of string theory shows that the perturbative vacuum is unstable. This instability implies that there exists a true vacuum to which a tachyon field $\varphi$ tends [53]. The foundations of this process are given by a model of effective field theory [54] with a Lagrangian of the DBI form
@@
This model belongs to a wider class of DBI theories described by the Lagrangian [55]
${\cal{L}}=-f(\theta)^{-1}\sqrt{1-f(\theta)g^{\mu\nu}\theta_{,\mu}\theta_{,\nu}}-V(\theta)\,.$ (3)
Making use of a field redefinition $\varphi=\int f(\theta)d\theta$, identifying $f(\theta)^{-1}=U(\varphi)$, and setting $V(\theta)=0$, yields a transformed Lagrangian of the form (2). A Lagrangian of the form (3) appears in the so-called brane inflation models [27, 56]. In these models, the motion of a D3-brane in a warped throat region of a compact space drives inflation, and the DBI field corresponds to the position of the D3-brane. In Ref. [27], a specific form of the warp factor $f$ is chosen to facilitate the phenomenological oscillating sound speed. In contrast, the tachyon potential $U$, corresponding to $f^{-1}$ in Eq. (3), is chosen to have an inflection point, as in the PLLS case, and thereby yields enhanced PBH production.
The advantage of the $k$-essence approach to PBH formation is due to its simple mechanism of ultra-slow-roll dynamics as a consequence of the inflection point. The models of this type yield straightforward predictions easily confronted with observations. Besides, as we show in the last section (see also Refs. [55]), the amplitude of the primordial non-Gaussianity in both PLLS and Tachyon models is well within observational constraints. More generally, the merit of the tachyon model considered here is in its fundamental origin in low-energy string theory and brane dynamics.
@@
We use the metric signature (+,,,)(+,-,-,-). Unless specified otherwise, we use the system of units in which $c=\hbar=1$. For convenience, we introduce a length scale $\ell$ so that $\ell\gg 1/M_{\rm Pl}$, where $M_{\rm Pl}=\sqrt{1/(8\pi G}$ is the reduced Planck mass. By $\mathcal{L}$ and $\mathcal{H}$, we denote the Lagrangian and Hamiltonian multiplied by $\ell^{4}$ so that our $\mathcal{L}$ and $\mathcal{H}$ are dimensionless. By $\varphi$ and $\phi$, we denote the scalar fields connected by $\varphi=\ell^{2}\phi$ and have the dimension of length and mass, respectively.
@@ 2 The PLLS Model
where
$X=g^{\mu\nu}\varphi_{,\mu}\varphi_{,\nu}$ (5)
is assumed positive. Here, $\varphi$ is a scalar field of dimension of length, and $\alpha\geq 1$ is a constant. The physical field $\phi$, related to $\varphi$ via
@@
[…]
${\cal H}(\eta^{\mu},\varphi)=\eta^{\mu}\varphi_{,\mu}-{\cal L}(\varphi_{,\mu},\varphi),$ (8)
with conjugate variables satisfying the conditions
$\varphi_{,\mu}=\frac{\partial{\cal H}}{\partial\eta^{\mu}},$ (9)
$\eta^{\mu}=\frac{\partial{\cal L}}{\partial\varphi_{,\mu}}.$ (10)
By making use of Eqs. (8), (10), and the Euler-Lagrange equation
$\left(\frac{\partial{\cal{L}}}{\partial\varphi_{,\mu}}\right)_{;\mu}=\frac{\partial{\cal{L}}}{\partial\varphi},$ (11)
we find
${\eta^{\mu}}_{;\mu}=-\frac{\partial{\cal H}}{\partial\varphi}.$ (12)
Equations (9) and (12) are the covariant Hamilton equations. The most general $k$-essence Lagrangian is a function of the form $\mathcal{L}=\mathcal{L}(X,\varphi)$, where $X$ is given by (5). From (10) it follows that the quantity
$\eta^{2}=g_{\mu\nu}\eta^{\mu}\eta^{\nu}=4X(\mathcal{L}_{X})^{2}$ (13)
is positive. Here, the subscript X denotes a derivative with respect to $X$. Then, one can introduce the four-velocity
$u_{\mu}=\frac{g_{\mu\nu}\eta^{\nu}}{\eta}=\epsilon\frac{\varphi_{,\mu}}{\sqrt{X}},$ (14)
where, $\epsilon$ is +1+1 or 1-1 according to whether $\varphi_{,0}$ is respectively positive or negative. This choice of $\epsilon$ guarantees that $u_{0}$ is always positive. The quantity $\eta$ in (14) is a square root of (13),
$\eta\equiv 2\epsilon\sqrt{X}{\cal L}_{X}=\epsilon\frac{\partial{\cal L}}{\partial\sqrt{X}}.$ (15)
We can regard the variable $\eta$ as a conjugate to $\sqrt{X}$ as described in [57]. Next, using (8) and (10) one can express $\mathcal{H}$ as a function of $X$ and $\varphi$,
${\cal H}=2X{\cal L}_{X}-{\cal L}.$ (16)
Then, from (13) it follows that $X$ is an implicit function of $\eta^{2}$ and $\varphi$ which we can (in principle) solve for $X$ to obtain $X=X(\eta^{2},\varphi)$. Hence, in a $k$-essence type of theory, one can regard the Hamiltonian as a function of $\eta^{2}=g_{\mu\nu}\eta^{\mu}\eta^{\nu}$ and $\varphi$.
Using the definitions of the velocity (14) and expansion rate
$H\equiv\frac{1}{3}{u^{\mu}}_{;\mu}\,$, (17)
the covariant Hamilton equations (9) and (12) can be expressed as
$\dot{\varphi}=\frac{\partial{\cal H}}{\partial\eta},$ (18)
$\dot{\eta}+3H\eta=-\frac{\partial{\cal H}}{\partial\varphi}.$ (19)
where a quantity $f$ with an over-dot means $\dot{f}\equiv u^{\mu}f_{,\mu}$. In the comoving frame, the over-dot becomes a derivative with respect to time, and the expansion rate becomes $H=\dot{a}/a$ as usual. For the PLLS Lagrangian (4) we find
@@
In this model, the variable $X$ can be expressed as an explicit function of $\eta$. By making use of (4) and (15) we find
$\frac{X}{2}=\left(\frac{\eta^{2}}{2\alpha^{2}}\right)^{1/(2\alpha-1)}.$ (21)
@@
The Hamilton equations (18) and (19) govern the background dynamics. Introducing the e-fold number
@@
Instead of the two first-order Hamilton equations, one can use the second-order field equation of motion. This equation can be obtained directly from the Lagrangian or combining Eqs. (28) and (29). Either way one finds [^1] A comparison with Ref. [31] reveals that the factor $\alpha(2\alpha-1)$ in the denominator of the last term on the left-hand side of Eq. (31) is missing in their equation (2.9).
@@ 3 The Tachyon model
As in the PLLS model, we choose the potential $U(\varphi)$ defined by (7). The parameters of the potential are $\ell=10^{6}M_{\rm Pl}^{-1}$, $n=3$, and $V_{0}=10^{-16}M_{\rm Pl}^{4}$ as before. For this model we choose a slightly different $\lambda=7.502\cdot 10^{-6}$. Regarding the initial values for $\varphi$ and $\eta$, we first choose $\phi_{\rm in}$ in units of $M_{\rm Pl}$ of the order of those in Ref. [31] with the corresponding $\varphi_{\rm in}=\ell^{2}\phi_{\rm in}$. Then, we tune $\varphi_{\rm in}$ and $\eta_{\rm in}$ to obtain plateau-like solutions as shown in Fig. 2.
Figure 3: The slow-roll parameters $\varepsilon_{1}$ (left panel) and $\varepsilon_{2}$ (right panel) versus NN in the Tachyon model for $\phi_{\rm in}=1.1M_{\rm Pl}$ (full line) and $1.0520M_{\rm Pl}$ (dashed line). The remaining corresponding parameters are listed in Table 1.
As in the PLLS case, the scalar field exhibits a plateau owing to the inflection point of the potential at $\phi=\phi_{0}$. As shown in Fig. 2, the extension of the plateau depends on $\phi_{0}$. In this flat region, slow-roll conditions will not be met, and inflation will enter a temporary ultra-slow-roll regime. In Fig. 3 we plot the slow-roll parameters as a function of $N$. During this period, as we will demonstrate in the next section, the non-constant mode of the curvature fluctuations, which would decay exponentially in the slow-roll regime, actually grows in the ultra-slow-roll regime. As a consequence, the curvature power spectrum will be enhanced at specific scales that can potentially collapse, forming primordial BHs in the early Universe.
@@ 4 Curvature perturbations
The power spectrum of curvature perturbations is obtained from the two-point correlation function
$\langle\hat{\zeta}_{q}\hat{\zeta}_{q^{\prime}}\rangle=(2\pi)^{3}\delta(\mbox{$q$}+\mbox{$q$}^{\prime})|\zeta_{q}|^{2},$ (49)
where $\hat{\zeta}_{q}$ is the operator associated with the curvature perturbation $\zeta$ [60]. We define
@@
where $\zeta_{q}(N_{q})$ is the solution to Eq. (47) taken at the point $N_{q}$ which for given $q$ satisfies (48). The curvature perturbation spectrum needs to satisfy the Harrison-Zeldovich spectrum near $q=q_{\rm CMB}=0.05$ Mpc-1
@@
Note that Eq. (53) follows from Eqs. (44)-(45) via (46) and (47). Since the derivation of Eqs. (44)-(45) does not involve the slow-roll assumption, the quantities $\varepsilon_{1}$ and $\varepsilon_{2}$ that appeear in Eqs. (53)-(55) are not necessarily small. In our numerical integration of Eq. (53) with (54)-(54), we will treat the slow-roll parameters $\varepsilon_{1}$ and $\varepsilon_{2}$ as exact background quantities, not subject to the slow-roll conditions.
@@
where the conformal time $\tau$ is defined by
$d\tau=\frac{dt}{a}=\frac{dN}{aH}.$ (61)
Hence, the initial values of $v_{q}$ and $v_{q}^{\prime}$ at $N_{{\rm in},q}$ are determined by (60) up to an arbitrary phase. The simplest choice is
@@ 4.2 Approximate spectrum
where for each $q$, the quantities $H$, $c_{\rm s}$, and $\varepsilon_{1}$ take on their horizon crossing values at the corresponding $N_{q}$. It is important to stress that in the models considered here the perturbations evolve through a phase which departs from the slow-roll regime, as is evident from Fig. 3. Nevertheless, we will use (67) for the sake of comparison. However, we expect a substantial departure of the power spectrum (67) from the exact one obtained by numerically integrating Eq. (53).
The approximate spectrum obtained in this way will not in general have the correct observed value at the pivot scale $q_{\rm CMB}$. To satisfy the proper normalization of $\mathcal{P}_{\rm S}(q)$ at $q=q_{\rm CMB}$, we can introduce a constant factor $\bar{c}_{0}$. To wit, we define
@@ 5 PBH formation
The quantity $\zeta(r)$ is spherically symmetric comoving curvature perturbation conserved on superhorizon scales [63]. The corresponding density contrast $\delta_{r}\equiv\delta\rho/\rho(r,t)$ is given by [9]
$\delta_{r}=-\frac{8}{9}\frac{1}{a^{2}H^{2}}e^{-5\zeta(r)/2}\nabla^{2}e^{\zeta(r)/2},$ (70)
Keeping the linear terms only, we find
$\delta_{r}\simeq-\frac{4}{9}\frac{1}{a^{2}H^{2}}\nabla^{2}\zeta(r).$ (71)
In the momentum space we obtain
@@
Hence, the two-point correlation function (49) dominates the power spectrum of the density contrast via the linear relation (72). We find
@@
In our analysis, we assume that both the density contrast $\delta$ and curvature perturbation $\zeta$ are approximately Gaussian variables. In principle, a non-Gaussian contribution to the statistics of the primordial curvature perturbations could modify the PBH abundance. We will provide a brief comment on non-Gaussianity at the end of this section.
As we will shortly see, the PBH mass fraction $\beta(M)$ will be mainly described by the smoothed variance $\sigma$ of the density contrast $\delta$, with rough functional dependence as
$\beta\propto\left(-\frac{\delta_{\rm c}}{2\sigma^{2}}\right),$ (74)
where $\delta_{\rm c}$ is the critical overdensity collapse threshold. Hence, the larger the variance $\sigma$, the larger the mass fraction $\beta$. We define the smoothed variance of $\delta$ over the smoothing scale $R=1/(aH)$
@@
where $W$ is a smoothing window function. It is not obvious what the best choice for the smoothing function is. There is a priori no reason to choose one smoothing over another [64, 65]. A common convenient choice is a Gaussian window function, although a top-hat smoothing function was occasionally used in the past [66, 67]. However, the top-hat window function is sensitive to scales well within the horizon [68] which requires careful treatment [69, 70].
Ando, Inomata, and Kawasaki [71] investigated the dependence of the variance $\sigma(R)$ on the choice between $x$-space top hat, gaussian, and $k$-space top hat window function (see also Refs. [72, 73]). In particular, for a simple scale invariant power spectrum $\mathcal{P}_{\rm S}=A_{\rm S}$, they found that the variances $\sigma(R)$ differ in proportions 1:0.29:0.22 for the $x$-space top hat, Gaussian, and $k$-space top hat window functions, respectively. We have also checked the window function dependence of $\sigma(R)$ assuming a gaussian shape power spectrum $\mathcal{P}_{\rm S}(q)\propto\exp[-R^{2}(q-R_{0}^{-1})^{2}/2]$ peaked at various $R_{0}^{-1}$. For example, for $R_{0}=R$, we have found the proportions 1:0.36:0.27 for the mentioned window functions, respectively. These proportions do not change significantly with varying $R_{0}$ in the interval 0.3 - 10, where they remain roughly 1:1/3:1/4.
Following Green et al. [68], we choose gaussian
$W(q,R)=e^{-q^{2}R^{2}/2},$ (77)
as a more conservative choice than the $x$-space top-hat window function.
@@ 5.1 PBH mass fraction
5.1.1 Press-Schechter (PS) formalism
In the Press-Schechter formalism of gravitational collapse, the mass fraction of PBHs of mass $M$ is described by the probability that the overdensity $\delta$ is above a certain threshold value $\delta_{c}$ for collapse. Assuming $\delta$ is a Gaussian random variable with mass (or scale) dependent variance, the mass fraction $\beta(M)$ at the time of formation is then given by [76, 77, 78]
@@ 5.1.1 Press-Schechter (PS) formalism
5.1.2 Critical collapse and peaks (CCP) formalism
@@
5.1.3 Comparison
The CCP formalism is generally considered to be more refined [65, 83] compared with the PS. Whereas the PS formalism has one universal condition that the density must be above the threshold value, the CCP formalism introduces an additional constraint: PBHs form at peaks of the density [64, 81]. In particular, as pointed out in Ref. [84], the wide variety of the threshold value implies the difficulty in the PS approach: the PBH formation criterion is not a simple matter of the universal threshold for a one-point value of some averaged random field, but we need to deal with the spatial dependence of the local overdensity profile. The two methods have been compared in a comprehensive recent analysis by Yoo et al. [85] (see also Pi et al. [86]) that takes into account the mutually related issues of PBH formation criterion, statistical treatment of non-linear variables, and use of a window function.
To compare these two methods quantitatively, we can use the toy-model spectrum (78). In this model, the function $\beta(M)$ may be approximated by a similar top-hat profile